55 shrimps = 32 oz.
so how many shrimps (x) = 3 oz

Answer by CharlesG2(828)

55 shrimps = 32 oz.
so how many shrimps (x) = 3 oz
55 shrimps / 32 oz. = x shrimps / 3 oz.
55*3=32*x cross multiplying
165=32*x
165/32=x
160/32 + 5/32 = x
5 + 5/32 = x
5.15625 shrimps = x